If 𝑓 is a continuous function on [0,1], show that limπ‘›β†’βˆž ∫ 𝑛𝑓(π‘₯) 1+𝑒2π‘₯ 2 1 0 𝑑π‘₯ = πœ‹ 2 𝑓(0).Β 


1
Expert's answer
2021-11-07T18:11:20-0500

Let


"I=\displaystyle{\int^{1}_{0}}\frac{nf(x)}{1+u^2x^2}dx=I_1+I_2"


where


"I_1=\displaystyle{\int^{n^{-1/3}}_{0}}\frac{nf(x)}{1+u^2x^2}dx"


and


"I_2=\displaystyle{\int^{1}_{n^{-1/3}}}\frac{nf(x)}{1+u^2x^2}dx"


"|f(x)|\u2264M" because f is continuous. We have:


"I_2\le\displaystyle{\int^{1}_{n^{-1/3}}}|\frac{nf(x)}{1+u^2x^2}|dx\le \displaystyle{\int^{1}_{n^{-1/3}}}\frac{nM}{1+u^2x^2}dx \le \displaystyle{\int^{1}_{n^{-1/3}}}\frac{nM}{1+u^2(n^{-1/3})^2}dx\\"

"=\frac{nM}{1+n^2(n^{-1/3})^2}(1-n^{-1/3})=o(1)"


Put


"I_3=\displaystyle{\int^{n^{-1/3}}_{0}}\frac{nf(0)}{1+n^2x^2}dx"


We have:


"|I_1-I_3|=|\displaystyle{\int^{n^{-1/3}}_{0}}\frac{n(f(x)-f(0))}{1+n^2x^2}dx|\le"


"\le \displaystyle{\sup_{n\in [0,n^{-1/3}]}}|f(x)-f(0)|\cdot \displaystyle{\int^{n^{-1/3}}_{0}}\frac{ndx}{1+n^2x^2}dx=o(1)\cdot \displaystyle{\int^{n^{-1/3}}_{0}}\frac{d(nx)}{1+n^2x^2}dx="


"=o(1)\cdot arctan(n\cdot u^{-1/3})=o(1)\cdot O(1)"


Hence

"I=I_1+I_2=I_3+(I_1-I_3)+I_2=I_3+o(1)+o(1)"


where

"I_3=f(0)\displaystyle{\int^{n^{-1/3}}_{0}}\frac{d(nx)}{1+u^2x^2}dx=f(0)\cdot arctan(n\cdot u^{-1/3})=f(0)\frac{\pi}{2}+o(1)"


So,


"I\to f(0)\frac{\pi}{2}" as "n\to \infin"